# FeDaL: Federated Dataset Learning for Time Series Foundation Models

### v1

@@ Abstract
Dataset-wise heterogeneity introduces significant domain biases that fundamentally degrade generalization on Time Series Foundation Models (TSFMs), yet this challenge remains underexplored. This paper rethink the development of TSFMs using the paradigm of federated learning. We propose a novel Federated Dataset Learning (FeDaL) approach to tackle heterogeneous time series by learning dataset-agnostic temporal representations. Specifically, the distributed architecture of federated learning is a nature solution to decompose heterogeneous TS datasets into shared generalized knowledge and preserved personalized knowledge. Moreover, based on the TSFM architecture, FeDaL explicitly mitigates both local and global biases by adding two complementary mechanisms: Domain Bias Elimination (DBE) and Global Bias Elimination (GBE). FeDaL’s cross-dataset generalization has been extensively evaluated in real-world datasets spanning eight tasks, including both representation learning and downstream time series analysis, against 54 baselines. We further analyze federated scaling behavior, showing how data volume, client count, and join rate affect model performance under decentralization.
@@ 1 Introduction
Time series data underpin critical decisions across domains such as climate bi2023accurate , traffic deng2024multi , and healthcare wang2024medformer , where tasks like forecasting lim2021time , imputation chen2024federated , classification zhang2025timesbertbertstylefoundationmodel , and anomaly detection liu2024timer are essential for operational intelligence. While early work emphasized task-specific models tailored to narrow objectives zhou2022fedformer , recent advances have shifted toward general-purpose time series foundation models (TSFMs) liu2024timer ; das2023decoder , inspired by the success of large language models. These TSFMs aim to generalize across various tasks and domains, but doing so requires access to both diverse public datasets and large-scale private data. Federated Learning (FL) mcmahan2017communication presents a scalable framework for training TSFMs collaboratively chen2024federated , as it enables training across distributed data while preserving privacy and reducing the computational burden associated with centralized training ye2024openfedllm .
Figure 1: Key domain biases in heterogeneous time series: temporal resolution bias, physical constraint bias, and pattern transition bias. Information Lost: Low-resolution sequences encode less detail within a fixed lookback window. Pattern Break: Changes at specific time points (or segments). FL Settings: FedAvg-Independent (each dataset as a client), FedAvg-Mixed (datasets mixed and split into two clients), Stand-alone (no aggregation). Other settings follow Section 4. Lower MSE indicates better performance. Best viewed in color and zoomed in.
@@
Despite its promise, training TSFMs via FL introduces unique challenges, as robust generalization requires addressing domain biases caused by heterogeneous datasets, with biases that are both diverse and underexplored in prior work chen2024federated . Unlike text/images, time series vary widely in resolution, physical constraints, and temporal dynamics, complicating the extraction of domain-invariant patterns. Although FL inherently supports heterogeneity through its decentralized nature, existing works chen2024federated primarily focus on aligning trends across domains, overlooking deeper, dataset-specific structural biases that can severely hinder generalization. We identify three key types of such bias (Figure˜1): (1) Temporal resolution bias, sequences with different sampling rates encode inconsistent contextual information under a fixed window (e.g., 120 steps cover five days in hourly weather data but only two hours in minute-level energy data); (2) Physical constraint bias, domain-specific dynamics governed by unrelated physical laws (e.g., temperature variation vs. electric current) reduce representational transferability; and (3) Pattern transition bias, initially similar trends across domains may diverge sharply due to exogenous events (e.g., traffic vs. web visits), breaking assumptions of shared temporal structures. These biases pose fundamental obstacles to learning robust and cross-domain transferable patterns via FL, ultimately limiting the generalization capacity of the learned TSFMs.
@@
This paper introduces Federated Dataset Learning (FeDaL), a brand-new framework for training Time Series Foundation Models under federated settings, with the goal of capturing generalized temporal patterns from heterogeneous datasets. To address domain-level heterogeneity, FeDaL incorporates two complementary components: Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), operating from the client and server perspectives, respectively. DBE approximates and disentangles client-specific biases in a context-agnostic manner by explicitly modeling and storing local bias representations. GBE enhances global model generalization by aligning client contributions through gradient-level correction and privacy-preserving, fine-grained server-side tuning, enabling adaptive bias mitigation during aggregation. Together, these components allow the pretrained TSFM to learn domain-invariant temporal representations, facilitating strong cross-domain generalization while maintaining privacy and scalability. Our key contributions are summarized as follows:
  • We identify three key domain biases (temporal resolution, physical constrains, and pattern transition) that hinder the learning of general temporal patterns in FL-based TSFM training.
@@
  • We propose Federated Dataset Learning (FeDaL), a novel framework for learning TSFMs in federated settings, which explicitly addresses dataset heterogeneity through context-agnostic Domain Bias Elimination (DBE) and server-side Global Bias Elimination (GBE).
@@
  • We perform the first systematic analysis of TSFM scaling behaviors from a federated learning perspective, offering empirical insights into how data volume, client population, and participation rate affect model generalization under decentralized settings.
  • Extensive experiments on real-world time series demonstrate that FeDaL enables strong in-domain and out-of-distribution generalization, consistently outperforming state-of-the-art baselines across forecasting, imputation, classification, and anomaly detection tasks.
@@ 2 Preliminaries
Pre-trained Time Series Models
Pretraining has emerged as a key strategy for building effective time series models, supporting both task-specific nie2022time ; jin2023time and general-purpose TSFMs goswami2024moment ; woo2024unifiedtraininguniversaltime ; shi2024time . Most existing TSFMs adopt masked reconstruction at the time-point zeng2023transformers or patch level liu2024timer ; garza2023timegpt , achieving strong zero-shot performance. Recent efforts such as Moment goswami2024moment , Moirai woo2024unifiedtraininguniversaltime , TimeFM das2024decoder , and Chronos ansari2024chronos push this further with large-scale masked pretraining to enhance downstream adaptation. However, these models are often tailored to forecasting and show limited generalization to tasks like classification or anomaly detection. Moreover, their reliance on centralized, large-scale datasets restricts real-world applicability, where data is often siloed and heterogeneous chen2024federated ; zhuang2023foundation . Federated learning mcmahan2017communication offers a practical alternative by enabling collaborative model training across decentralized domains while preserving privacy chen2024federated . Building on this paradigm, we propose a brand-new federated framework for robust TSFM pretraining under decentralized and domain-heterogeneous settings, enabling broad adaptation across downstream tasks and domains.
Federated Learning in Time Series
FL provides a practical framework for collaborative training without centralized data fusion mcmahan2017communication , especially important given rising privacy concerns and the emergence of foundation models zhuang2023foundation . In time series analysis, FL has enabled both personalized models liu2024time and, more recently, generalized TSFMs chen2024federated , drawing increasing attention chen2023prompt ; chen2023federated ; sun2024hifi . While existing methods focus on client-level heterogeneity, they often overlook domain-level bias—a key challenge in time series where clients frequently hold non-overlapping, domain-specific data. These domain biases result in inconsistent optimization paths and poorly aligned latent patterns, limiting TSFM generalization. To address this, we introduce FeDaL, a brand-new FL paradigm that explicitly models and aligns domain-specific and global representations. By mitigating domain bias, FeDaL enables the training of TSFMs with improved cross-domain generalization.
Problem Definition
@@
We consider a FL setting with a server and $N$ clients, where each client $i$ owns a local time series dataset ${\mathcal{D}}_{i}$ drawn from one or more domains. These datasets are inherently heterogeneous due to variations in sampling rates, domain-specific characteristics, and temporal dynamics. The objective is to collaboratively train a unified TSFM that generalizes across clients and supports diverse downstream tasks. Formally, we aim to minimize the global objective: $F(\theta):=\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}\frac{n_{i}}{n}F_{i}(\theta_{i};{\mathcal{D}}_{i})$, where $F_{i}(\cdot)$ denotes the local loss, and $_{i}, nn$ are the local and total sample counts, respectively. This work asks: Can we train a TSFM under federated constraints that captures cross-domain temporal structure while minimizing the impact of domain-specific biases?
3 Methodology: Federated Dataset Learning
Overview
Heterogeneous time series datasets often introduce domain biases, causing local models to overfit to spurious patterns tied to particular datasets or clients. When these locally biased models are aggregated, their misaligned representations and divergent optimization trajectories result in a globally biased TSFM with limited generalization. These challenges are especially pronounced in massive, high-dimensional, and structurally diverse time series data. To address this, we propose two complementary mechanisms—Domain Bias Elimination (DBE) and Global Bias Elimination (GBE)—that operate at the client and server levels, respectively. Together, they explicitly mitigate domain-level bias and enable the alignment of complex temporal patterns into a unified, generalizable representation. The detailed algorithmic implementation of our FeDaL (Left, Figure˜2) is provided in Algorithm˜1. The remainder of this section details the designs and roles of DBE and GBE.
Figure 2: (Left) FedDaL with two clients example: DBE alleviates domain bias from the client perspective, while GBE addresses global bias from the server side. (Right) Each client employs a Decoder-only Transformer liu2024time ; liu2024timer , extracting trend and season biases from latent representations via plug-and-play DBE block.
@@ 3.1 Domain Bias Elimination
When client models overfit to domain-specific artifacts, their latent representations tend to entangle general temporal features with spurious local biases, undermining the quality of representation learning and hindering downstream generalization. To mitigate this, we introduce Domain Bias Elimination (DBE) (refer to  Right, Figure˜2). DBE decomposes the latent representation of masked input patches into context-agnostic components (trend and seasonal), from which a trainable local bias vector is estimated. This vector is injected back during reconstruction to isolate and suppress domain-specific biases. Additionally, we regularize its alignment to a global bias reference, guiding the model toward disentangling spurious local patterns from generalizable temporal structure.
Given a masked input sequence ${\bm{X}}_{i}^{\text{masked}}$ on client $i$, its latent representation is obtained using the local backbone $f_{\theta^{b}_{i}}(\cdot)$. We introduce a plug-and-play module, the DBE block, which incorporates time series decomposition wu2021autoformer and averaging to explicitly estimate local bias. Specifically, we first decompose the masked input representations into trend and seasonal components, as follow:
$\mathbf{h}_{i}^{\text{trend}},\mathbf{h}_{i}^{\text{season}}=\texttt{Decomp}(f_{\theta^{b}}({\bm{X}}_{i}^{\text{masked}}),\tau),\quad\mathbf{b}_{i}^{\text{season}},\mathbf{b}_{i}^{\text{trend}}=\texttt{Mean}(\mathbf{h}_{i}^{\text{season}}),\texttt{Mean}(\mathbf{h}_{i}^{\text{trend}})$ (1)
where $\tau$ denotes the decomposition granularity. The components $\mathbf{h}_{i}^{\text{trend}},\mathbf{h}_{i}^{\text{season}}$ capture generalizable, context-independent temporal patterns. By averaging these components, we extract persistent directional deviations in the latent space zhang2023eliminating , which serve as compact and effective approximations of client-specific biases arising from the accumulation of complex time series dataset heterogeneity. We estimate the local bias vector by summing the two components: $\hat{\mathbf{b}}_{i}^{p}=\mathbf{b}^{\text{trend}}_{i}+\mathbf{b}^{\text{season}}_{i}$. Both $\mathbf{b}_{i}^{\text{trend}}$ and $\mathbf{b}_{i}^{\text{season}}$ are kept trainable to adaptively capture evolving client-specific characteristics. During training, the model reconstructs the original input by injecting the bias vector into the latent space:
${\mathcal{L}}_{i}(\theta_{i}):=\mathbb{E}\left[\left\|f_{\theta^{h}}(f_{\theta^{b}}({\bm{X}}^{\text{masked}})+\hat{\mathbf{b}}_{i}^{p})-{\bm{X}}^{\text{orig}}\right\|^{2}\right].$ (2)
Here, $\theta_{i}=[\theta^{b},\theta^{h},\hat{\mathbf{b}}_{i}^{p}]$ includes all local learnable parameters, and ${\bm{X}}_{i}^{\text{orig}}$ is the original input sequence. To promote alignment across clients and reduce divergence caused by heterogeneous biases, we introduce an explicit regularization term based on bias approximation to rewrite Eq. 2 as:
${\mathcal{L}}_{i}(\theta_{i}):=\mathbb{E}[\|f_{\theta^{h}}(f_{\theta^{b}}({\bm{X}}^{\text{masked}})+\hat{\mathbf{b}}_{i}^{p})-{\bm{X}}^{\text{orig}}\|^{2}]+\lambda\|\hat{\mathbf{b}}_{i}^{p}-\mathbf{b}^{g}\|^{2}.$ (3)
This encourages disentanglement of biased and unbiased components in the local representation, guiding the projection head to emphasize generalizable features. The global bias vector $\mathbf{b}^{g}$ is computed once during initialization (refer to Line 3 in Algorithm 1), and alignment is enforced via a mean squared error (MSE) penalty between the local and global biases. The regularization strength is controlled by a tunable coefficient $\lambda$, optimized jointly via SGD. Since the bias term in Eq. 3 depends on aggregated statistics across the full local dataset—while SGD operates on mini-batches—we approximate the global signal using an exponential moving average zhang2015deep across batches:
$\mathbf{b}_{i}^{\text{trend}}=(1-\mu)\mathbf{b}_{i,\text{old}}^{\text{trend}}+\mu\mathbf{b}_{i,\text{new}}^{\text{trend}},\quad\mathbf{b}_{i}^{\text{season}}=(1-\mu)\mathbf{b}_{i,\text{old}}^{\text{season}}+\mu\mathbf{b}_{i,\text{new}}^{\text{season}},$ (4)
where $\mu$ is a smoothing coefficient, and “new“ and “old“ denote current and previous batch-wise estimates, respectively. This allows stable bias tracking throughout local training.

### v2

# FeDaL: Federated Dataset Learning for General Time Series Foundation Models

@@ Abstract
Dataset-level heterogeneity introduces significant domain biases that fundamentally degrade generalization on general Time Series Foundation Models (TSFMs), yet this challenge remains underexplored. This paper rethinks the from-scratch training of TSFMs using the paradigm of federated learning. We propose a novel Federated Dataset Learning (FeDaL) approach to tackle heterogeneous time series by learning dataset-agnostic temporal representations. Specifically, the distributed architecture of federated learning is a nature solution to decompose heterogeneous TS datasets into shared generalized knowledge and preserved personalized knowledge. Moreover, based on the TSFM architecture, FeDaL explicitly mitigates both local and global biases by adding two complementary mechanisms: Domain Bias Elimination (DBE) and Global Bias Elimination (GBE). FeDaL‘s cross-dataset generalization has been extensively evaluated in real-world datasets spanning eight tasks (including various regression and classification), against 54 baselines. We further analyze federated scaling behavior, showing how data volume, client count, and join rate affect model performance under decentralization. Our code is publicly available at https://github.com/shengchaochen82/FeDaL.
@@ 1 Introduction
Time series analysis plays critical roles in decision systems such as weather forecasting (Bi et al., 2023), medical symptom classification (Wang et al., 2024c), industrial anomaly detection (Xu et al., 2022), and imputing missing data in wearable sensors (Wu et al., 2020). Recent advances follow two main directions. Time Series Foundation Models (TSFMs) pretrain on large-scale, multi-domain datasets to capture transferable representations, achieving strong zero-shot or few-shot performance but remaining largely task-specific, for example in forecasting (Liu et al., 2024c; Shi et al., 2024; Ansari et al., 2024; Das et al., 2023) or classification (Feofanov et al., 2025). In contrast, Time Series Pattern Machines (TSPMs) pursue architecture-level unification across tasks (Zhou et al., 2023; Wu et al., 2022a; Wang et al., 2024a). However, TSPMs are typically trained per dataset, restricting zero-shot generalization across unseen domains. Moreover, similar to TSFMs, they often assume centralized access to diverse datasets, which is unrealistic in practice since time series data are usually siloed and heterogeneous (Chen et al., 2025). This gap motivates the need for a general TSFM that combines architecture-level flexibility with data-level generalization, without centralizing data.
Federated FMs (FFMs) (Zhuang et al., 2023) provide a promising decentralized paradigm for training FMs by leveraging both public and private datasets. This strategy mitigates privacy concerns and alleviates the computational burden of centralized training, and has been extended to train general-purpose models by treating each domain or dataset as an independent client. Despite this promise, existing works (Chen et al., 2025) face two major challenges. First, they adopt a coarse-grained treatment of heterogeneity, focusing only on broad domain-level shifts (e.g., climate vs. healthcare) while overlooking dataset-specific structural biases within each domain. Consequently, the lack of robust cross-domain invariance limits scalability, as downstream tasks still require fine-grained tuning to adapt to individual datasets, preventing these models from functioning as true FMs.
@@
To make this concrete, we illustrate three representative dataset-level biases in Figure 1: (1) Temporal resolution bias, where sequences with different sampling rates encode inconsistent contexts under a fixed window (e.g., 120 steps cover five days of hourly weather but only two hours of minute-level energy data); (2) Physical constraint bias, where unrelated physical laws (e.g., temperature variation vs. electric current) reduce cross-domain transferability; and (3) Pattern transition bias, where initially similar trends diverge due to exogenous events (e.g., traffic vs. website visits), breaking assumptions of shared temporal structures. These biases illustrate that heterogeneity arises at both the client and global levels: temporal resolution and physical constraint biases primarily distort client-specific representations, while pattern transition bias amplifies during global aggregation.
Figure 1: Dataset-level biases. Information Lost: Low-resolution sequences capture fewer details under a fixed window. Pattern Break: Abrupt structural changes across time. Settings: FedAvg-Independent (each dataset as a client), FedAvg-Mixed (datasets pooled then split), and Stand-alone (no aggregation). ‘Transformer‘ means that results from the local Transformer model. Setup follow Sec. 4. Lower MSE is better.
@@
This paper introduces Federated Dataset Learning (FeDaL), a federated framework for training general TSFMs from scratch, designed to capture dataset-level patterns from heterogeneous datasets. FeDaL moves beyond coarse domain-level treatment to address dataset-level heterogeneity stemming from differences in sampling resolution, physical constraints, and transition patterns. Specifically, FeDaL integrates two complementary components: Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), operating at the client and server levels respectively. DBE approximates and disentangles client-specific biases in a context-agnostic manner by modeling local bias representations. GBE improves global generalization by aligning client updates via gradient-level correction and applying fine-grained server-side tuning to mitigate residual biases during aggregation. DBE and GBE enable FeDaL to produce TSFMs with domain-invariant representations that generalize across diverse datasets while preserving privacy and scalability. Our contributions are:
  • 1.
@@
    We propose Federated Dataset Learning (FeDaL), a FL framework for training general TSFMs from scratch, explicitly addressing dataset heterogeneity through context-agnostic Domain Bias Elimination (DBE) on clients and Global Bias Elimination (GBE) at server.
  • 2.
@@
    We provide the first systematic study of TSFM pretraining scaling behaviors across various tasks under federated learning, offering empirical insights into how data volume, client population, and participation rate shape model generalization in decentralized settings.
  • 3.
    Extensive experiments show that FeDaL-trained models outperforms advanced baselines across forecasting, imputation, anomaly detection, and classification, while achieving highly competitive performance against centralized TSFMs with far fewer parameters.
@@ 2 Preliminaries
Pre-trained Time Series Models.
Pretraining has become central to time series modeling, driving both task-specific models (Nie et al., 2022; Jin et al., 2023) and TSFMs (Goswami et al., 2024; Woo et al., 2024; Shi et al., 2024). Early efforts typically rely on masked reconstruction at the time-point (Zeng et al., 2023) or patch level (Liu et al., 2024c; Garza and Mergenthaler-Canseco, 2023), achieving strong zero-shot performance. More recent models such as Time-MoE (Shi et al., 2024), Moment (Goswami et al., 2024), Moirai (Woo et al., 2024), TimeFM (Das et al., 2024), and Chronos (Ansari et al., 2024) replace reconstruction with long-sequence prediction, further boosting zero-shot forecasting. However, these models remain task-oriented rather than universally applicable, unlike TSPMs that pursue architecture-level generality. In addition, they assume centralized access to massive datasets, an unrealistic requirement since time series are often siloed and heterogeneous (Chen et al., 2025; Zhuang et al., 2023). Federated learning (FL) (McMahan et al., 2017) offers a practical alternative by enabling decentralized training with privacy preservation.
Federated Foundation Models.
FFMs have emerged as a promising direction for scaling FMs without centralizing data (Zhuang et al., 2023). In TS, research on FFMs has begun to emerge. Time-FFM (Liu et al., 2024a) and PeFAD (Xu et al., 2022) leverages pretrained LLMs and applies lightweight tuning to achieve personalized predictions. In contrast to achieve multiple customized models, FFTS (Chen et al., 2025) takes a different approach by pretraining a TSFM on cross-domain datasets and then performing task-specific adaptation on downstream. It reduces local-global discrepancies to mitigate domain-wise heterogeneity. However, FFTS still falls short of being a true FM, since it don‘t support zero-shot inference and considers heterogeneity only at a coarse level, leaving dataset-level biases largely unresolved. To bridge this gap, we propose FeDaL, an FL framework TSFM training framework that explicitly addresses dataset-level heterogeneity and enables training a general TSFM from scratch with multiple downstream TS analysis task support.
@@
Problem Definition. We consider an FFM setting with a server and $N$ clients, where each client $i$ holds a local time series dataset ${\mathcal{D}}_{i}$. These datasets are heterogeneous due to differences in sampling resolution, physical constraints, and temporal dynamics, which induce dataset-specific biases that may not transfer across clients. The objective is to collaboratively train a general TSFM that generalizes that supports diverse downstream tasks. Formally, the global objective is $F(\theta)=\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}\frac{n_{i}}{n}F_{i}(\theta_{i};{\mathcal{D}}_{i})$, where $F_{i}(\cdot)$ denotes the local loss, and $_{i}, nn$ are the local and total sample counts. Our key question is: Can train a general TSFM from scratch under FL constraints that captures dataset-invariant temporal pattern while mitigating dataset-induced biases?
3 Methodology
Heterogeneous datasets introduce dataset-level biases, causing local models to overfit client-specific patterns that are not transferable across domains. For example, energy data may encode biases from grid cycles, while health data may reflect hospital-specific sampling protocols; such patterns are valid locally but break down in cross-domain settings. When aggregated, this misalignment produces a globally biased TSFM with limited generalization, a problem further amplified in large and structurally diverse datasets. To address this, we propose FeDaL (Figure 2) with two complementary mechanisms: Domain Bias Elimination (DBE) on clients and Global Bias Elimination (GBE) on server, where “domain“ is dataset-level heterogeneity rather than broad application classes. FeDaL mitigate dataset-induced biases and align temporal patterns into a unified, transferable space.
Figure 2: FeDaL: DBE reduces dataset-induced biases on the client side, whereas GBE improves alignment on the server. Locally, the plug-and-play DBE block (instead of DBE mechanism) captures trend and seasonal biases from latent representations.
@@ 3.1 Domain Bias Elimination
DBE (as shown in Figure 2) addresses dataset-induced biases by decomposing latent representations of masked input patches into trend and seasonal components, from which a trainable bias vector is derived. This bias vector is injected back during reconstruction to suppress dataset-specific artifacts (see Right, Figure 2), while a regularization term aligns it with a global bias reference, encouraging the model to disentangle non-transferable local patterns from generalizable temporal structures.
Following (Nie et al., 2022), we adopt an unsupervised patch reconstruction for local updating. For clarity, the client index $i$ is omitted in the following derivations. Given a masked sequence $\tilde{{\bm{X}}}$ obtained by patch-wise masking from original sequence $\bm{X}$}, its latent representation is obtained via the local backbone $f_{\theta^{b}}(\cdot)$. We introduce a plug-and-play DBE block to explicitly approximate dataset-induced biases. Instead of directly averaging latent features, we apply decomposition (Wu et al., 2021) to split them into trend and seasonal components to capture context-independent patterns as:
$\mathbf{h}_{i,t},\mathbf{h}_{i,s}=\operatorname{TimeDecomp}(f_{\theta^{b}}(\tilde{{\bm{X}}}),\tau),$ (1)
where $\tau$ is the decomposition granularity, $\mathbf{h}_{t}$ denotes the trend component, and $\mathbf{h}_{s}$ the seasonal component. We then average each component to extract persistent deviations from compact representations, producing bias vectors that capture stable dataset-specific shifts. This process as:
$\mathbf{b}_{t}=\operatorname{Mean}(\mathbf{h}_{t})\odot\bm{\gamma}_{t},\quad\mathbf{b}_{s}=\operatorname{Mean}(\mathbf{h}_{s})\odot\bm{\gamma}_{s}.$ (2)
Then we define the local bias vector $\mathbf{b}=\mathbf{b}_{t}+\mathbf{b}_{s}$. Compared with a plain mean of features, decomposition introduces an inductive bias: $\mathbf{b}_{t}$ captures low-frequency shifts, while $\mathbf{b}_{s}$ captures higher-frequency periodicities. Both are learnable, where the element-wise scaling factors $\bm{\gamma}_{t}$ and $\bm{\gamma}_{s}$ enable adaptation to evolving dataset patterns. To encourage disentanglement, the model reconstructs the input by injecting $\mathbf{b}$ into the latent features, the process can be formulated as:
${\mathcal{L}}(\theta)=\mathbb{E}[\|f_{\theta_{h}}(f_{\theta_{b}}(\tilde{{\bm{X}}})+\mathbf{b})-{\bm{X}}\|^{2}],$ (3)
where $\theta=[\theta_{b},\theta_{h},\mathbf{b}]$. This design ensures that dataset-specific deviations are absorbed by $\mathbf{b}$, forcing $f_{\theta_{b}}$ to focus on transferable structure. Directly reconstructing $\mathbf{b}$ alone would not constrain the backbone, whereas this additive formulation encourages disentanglement between invariant features and dataset-bound biases. To align local biases across clients, we rewrite local optimization objective Eq. 3 via introducing an explicit penalty with controlling coefficient $\lambda$ as:
${\mathcal{L}}(\theta)=\mathbb{E}[\|f_{\theta_{h}}(f_{\theta_{b}}(\tilde{{\bm{X}}})+\mathbf{b})-{\bm{X}}\|^{2}]+\lambda\|\mathbf{b}-\mathbf{b}^{g}\|^{2},$ (4)
Here $\mathbf{b}^{g}$ is initialized from aggregated client statistics and serves as a dynamic global prior, preventing drift. Although the penalty is applied to $\mathbf{b}$, its effect propagates to the backbone by constraining the separation of spurious and transferable patterns. Since bias depend on full-dataset statistics but training proceeds on mini-batches, we adopt an exponential moving average (EMA) (Zhang et al., 2015) to stabilize updates by gradually blending the old estimates with new batch-wise estimates:
$\mathbf{b}_{t}\!\leftarrow\!(1-\mu)\mathbf{b}_{t}^{\text{old}}+\mu\mathbf{b}_{t}^{\text{new}},\quad\mathbf{b}_{s}\!\leftarrow\!(1-\mu)\mathbf{b}_{s}^{\text{old}}+\mu\mathbf{b}_{s}^{\text{new}}.$ (5)
where $\mu$ is a smoothing factor. DBE can explicitly approximate dataset-specific biases, disentangle them from generalizable temporal patterns, and reduce inter-client divergence during aggregation.